# Networked Multiagent Reinforcement Learning for Peer-to-Peer Energy Trading

### v2

@@ Introduction
             Consensus step: $\omega_{t+1}^{i}\leftarrow\sum_{j=1}^{I}\widetilde{\omega}_{t}^{j}$.
@@
For each agent $i$, the action-value function is approximated by linear functions, i.e., $\hat{Q}(s,a;\omega)=$ $\omega^{\top}\phi(s,a)$ where $\phi(s,a)=\left[\phi_{1}(s,a),\cdots,\phi_{K}(s,a)\right]^{\top}\in\mathbb{R}^{K}$ is the feature associated with $(s,a)$. The feature function, $\phi_{k}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, for $k=1,\cdots,K$, is bounded for any $s\in\mathcal{S},a\in\mathcal{A}$. Furthermore, the feature functions $\left\{\phi_{k}\right\}_{k=1}^{K}$ are linearly independent, and for any $u\in\mathbb{R}^{K}$ and $u\neq 0,u^{\top}\phi$ is not a constant function over $\mathcal{S}\times\mathcal{A}$.
@@
The general convergence result of the consensus algorithm is proved in [1] under six assumptions. Our Assumption 1, 2, 3 and 4 directly correspond to four assumptions in [1]. The fifth assumption in [1] regarding agents’ time-varying neighborhood is trivially true here since we assume agents can communicate through the whole network, and the network topology does not change over time. Hence, the only remaining item to show is the uniform boundedness of agents’ reward $r_{i,t}$. By the definition of the market clearing price $P_{t}$ in (III), it is bounded between $FIT$ and $UR$. Together with the compactness assumption of the state space $\mathcal{S}_{i}$ (which is reasonable since an agent’s energy demand and battery/PV capacities are all bounded), the $R_{i}^{m}$ component in the reward function, as defined in (4), is uniformly bounded. The other term, the voltage violation penalty $R_{i}^{v}$, is uniformly upper bounded by 0 and lower bounded by $-\lambda NM$. Hence, $r_{i,t}$ is uniformly bounded for all $i$ and $t$, and the convergence results follow directly from Theorem 2 and 3 in [1]. ∎
@@
Remark 4. While the theorem establishes the convergence to a steady state of the critic and actor steps, it does not address how good the system reward is upon convergence. A key concern in this context is the potential inadequacy of the linear approximation of the critic function $\hat{Q}(s,a;\omega)$ in Assumption 3. In our model, the distribution network encompasses a multitude of agents, each exhibiting non-linear behaviors and interactions that may not be adequately captured by simple linear models. One natural idea is to deep neural networks (DNNs) instead, as they excel in capturing and modeling non-linear relationships. The downside of using DNNs, however, is that the proof of Theorem IV.1 is no longer applicable. Consequently, the convergence of the critic and actor steps remains an unresolved issue, meriting further investigation in future research.
Simulation Results
Test Data
@@
We test the algorithm frameworks using the IEEE 13-bus test feeder [25]. In our simulations, we set each bus to have one prosumer, excluding the substation bus; hence, there are 12 prosumers (and no pure consumers). The length of each time step is set to be 1 hour. The distribution network’s voltage limit is set to $[0.96\ pu,1.04\ pu]$. The amplifying parameter for voltage violation, $\lambda$, is set at $10^{4}$. The price ceiling and floor, that is, the utility rate $UR$ and feed-in tariff $FIT$, are set at 14 ¢/kWh and 5 ¢/kWh, respectively. The capacity of each agent’s PV is set to be 30kW, and the energy storage capacity is set to be 50kWh, with the charging and discharging efficiency being 0.95 and 0.9 for each agent. The PV generation of the prosumers and the total base load of the pure consumers on each bus have fixed diurnal shapes. These shapes represent the average values corresponding to each hour of the day. For a detailed explanation and due to page limitations her, please refer to our earlier paper [2]. Additionally, [2] provides more information on how the random PV output for each agent is generated. The power flow equations (7) are solved by the open-source distribution system simulator OpenDSS [26].
Deep Neural Network Implementation
@@
Comaprison of Three Different MARL Algorithms
@@
$\widehat{\nabla_{\theta_{i}}J_{i}\left(\theta_{i}\right)}=$ (30)
$\ \mathbb{E}_{s\sim\rho^{\theta},a_{i}\sim\pi_{\theta_{i}}}\left[\nabla_{\theta_{i}}\log\pi_{\theta_{i}}\left(a_{i}\mid s_{i}\right)Q_{i}^{\pi}\left(\bm{s};\bm{a_{1},\ldots,a_{I}}\right)\right],$
@@
While the action-value function $Q_{i}^{\pi}$ is the same as in (30), the key difference lies in the policy function. As stated in Section IV-A and reflected in (32), each agent $i$’s policy depends on all agents’ states $\bm{s}$. The other difference between the consensus MARL and the MADDPG algorithm is how the action-value function $Q_{i}^{\pi}$ is updated. In MADDPG, the function is updated through centralized training by a central authority; while in consensus MARL, updates are executed via a decentralized consensus process, as described in Section IV-B.
Numerical Results
Figure 3 displays the convergence curves of the 30-episode moving average of the episodic total reward in the distribution network for the three different MARL algorithms.
@@
It is evident that the mean episodic total reward gradually converges to a high level at the end of the training in all three frameworks. The results show that the consensus MARL algorithm outperforms the fully decentralized framework, which is in line with expectations, since the consensus-update framework benefits from improved approximation of the action-value function through communication. The better performance of consensus MARL over MADDPG, on the other hand, is surprising. Conventionally, one might anticipate a centralized algorithm to outperform decentralized algorithms. The explanation here can be traced to MADDPG’s decentralized execution. In MADDPG, each agent independently updates its actions based on its own local state, despite the centralized nature of the learning process. This approach of relying solely on local information appears to be less effective compared to the consensus update mechanism.
Figure 4 presents the market prices of the P2P market in chronological order over the last three days after thousands of training episodes.
@@
Notably, the market price in the consensus-update framework demonstrates a tendency to be higher during off-peak hours and lower during peak hours. This indicates that agents employing the consensus-update algorithm have developed more sophisticated trading strategies than the other two MARL frameworks. Specifically, the consensus-algorithm-induced strategies involve buying more and selling less during off-peak hours, and conversely, selling more and buying less during peak hours, when compared with the strategies produced by the purely decentralized algorithm.
Figure 5 shows the gradual reduction of the system voltage deviation after training. Initially, it is observed that voltage violations are significantly high across all three frameworks. However, the consensus framework demonstrates a rapid decline in these violations, converging swiftly and effectively to zero at the fastest rate. The remaining two frameworks demonstrate a slower rate of convergence, with the fully decentralized approach seemingly unable to fully eliminate voltage violations.
@@
Conclusion and Future Research
This study has developed a market and algorithmic framework to enable consumers’ and prosumers’ participation in local P2P energy trading. Utilizing reinforcement learning algorithms, we have automated bidding for agents while ensuring decentralized decision-making. The SDR-based market clearing addresses challenges with zero-marginal-cost resources and simplifies bidding. Additionally, our MARL framework includes voltage constraints of the physical network, setting a foundation for real-world applications.
@@
Additionally, while integrating transmission constraints into the current framework is straightforward (that is, they can be added to the power flow equations (7)), the allocation of transmission losses raises important equity considerations that must be addressed.

### v3

@@ 4.2 Consensus-update Actor-critic Algroithm
             Consensus step: $\omega_{t+1}^{i}\leftarrow\frac{1}{I}\sum_{j=1}^{I}\widetilde{\omega}_{t}^{j}$.
@@ 4.3 Convergence Results
For each agent $i$, the action-value function is approximated by linear functions, i.e., $\hat{Q}(s,a;\omega)=$ $\omega^{\top}\phi(s,a)$ where
$\phi(s,a)=\left[\phi_{1}(s,a),\cdots,\phi_{K}(s,a)\right]^{\top}\in\mathbb{R}^{K}$
is the feature associated with $(s,a)$. The feature function, $\phi_{k}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, for $k=1,\cdots,K$, is bounded for any $s\in\mathcal{S},a\in\mathcal{A}$. Furthermore, the feature functions $\left\{\phi_{k}\right\}_{k=1}^{K}$ are linearly independent, and for any $u\in\mathbb{R}^{K}$ and $u\neq 0,u^{\top}\phi$ is not a constant function over $\mathcal{S}\times\mathcal{A}$.
@@
The general convergence result of the consensus algorithm is proved in [1] under six assumptions. Our Assumption 1, 2, 3 and 4 directly correspond to four assumptions in [1]. The fifth assumption in [1] regarding agents’ time-varying neighborhood is trivially true here since we assume agents can communicate through the whole network, and the network topology does not change over time. Hence, the only remaining item to show is the uniform boundedness of agents’ reward $r_{i,t}$. By the definition of the market clearing price $P_{t}$ in (2), it is bounded between $FIT$ and $UR$. Together with the compactness assumption of the state space $\mathcal{S}_{i}$ (which is reasonable since an agent’s energy demand and battery/PV capacities are all bounded), the $R_{i}^{m}$ component in the reward function, as defined in (4), is uniformly bounded. The other two terms, the voltage violation penalty $R_{i}^{v}$, is uniformly upper bounded by 0 and lower bounded by $-\lambda NM$, and the capacity violation penalty $R_{i}^{c}$, is uniformly upper bounded by 0 and lower bounded by $-\lambda NC$. Hence, $r_{i,t}$ is uniformly bounded for all $i$ and $t$, and the convergence results follow directly from Theorem 2 and 3 in [1]. ∎
@@
Remark 4. While the theorem establishes the convergence to a steady state of the critic and actor steps, it does not address how good the system reward is upon convergence. A key concern in this context is the potential inadequacy of the linear approximation of the critic function $\hat{Q}(s,a;\omega)$ in Assumption 3. In our model, the distribution network encompasses a multitude of agents, each exhibiting non-linear behaviors and interactions that may not be adequately captured by simple linear models. One natural idea is to deep neural networks (DNNs) instead, as they excel in capturing and modeling non-linear relationships. The downside of using DNNs, however, is that the proof of Theorem 4.1 is no longer applicable. Consequently, the convergence of the critic and actor steps remains an unresolved issue, meriting further investigation in future research.
Remark 5. To guarantee convergence, all agents must use the consensus algorithm. It has been shown in [26] that this consensus-based MARL algorithm may lack robustness; specifically, a single adversarial attacker could disrupt the joint reward optimization process. A more resilient approach is proposed in the follow-up work [31], which could potentially be adapted to enhance the robustness of our framework. While the focus is not necessarily on adversarial agents, it would be valuable to address scenarios where some agents may use different algorithms or act randomly without using any algorithm at all. Exploring such robustness would be a promising future research direction.
Another challenge arises from the dynamic nature of multi-agent environments, where agents frequently enter or leave the system, resulting in an ever-changing environment. The convergence of consensus MARL depends on the exchange of information between neighboring agents (while our results address the exchange of information with all agents in the network, the algorithm is also effective for local neighborhood exchanges, as demonstrated in [1]). As the composition of agents evolves, the joint state and action spaces of neighboring agents can change, potentially impacting the learned policies. Transfer learning [32], which facilitates the transfer of knowledge between domains, may provide a viable solution to adapt policies in such dynamic settings. Specifically, techniques like policy distillation and policy reuse can help preserve useful behaviors while adapting to new conditions. Investigating the application of these techniques in the MARL context offers a promising direction for future research.
5 Simulation Results
5.1 Test Data
We test the algorithm frameworks using the IEEE 13-bus test feeder [33], whose topology is shown in Figure 5.
Figure 5: IEEE-13 test feeder
@@ 5.1 Test Data
In our simulations, we set each bus to have one prosumer, excluding the substation bus; hence, there are 12 prosumers (and no pure consumers). The length of each time step is set to be 1 hour. The distribution network’s voltage limit is set to $[0.98\ pu,1.01\ pu]$. The amplifying parameter for voltage violation, $\lambda$, is set at $10^{4}$. As for the line capacity constraint, we set an apparent power constraint of 15 kVA on the line between Node 684 and 671, and the amplifying parameter for capacity violation, $\gamma$, as $10^{2}$. The price ceiling and floor, that is, the utility rate $UR$ and feed-in tariff $FIT$, are set at 14 ¢/kWh and 5 ¢/kWh, respectively. The capacity of each agent’s PV is set to be 20kW, and the energy storage capacity is set to be 50kWh, with the charging and discharging efficiency being 0.95 and 0.9 for each agent. The PV generation of the prosumers and the total base load of the pure consumers on each bus have fixed diurnal shapes. These shapes represent the average values corresponding to each hour of the day. For a detailed explanation and due to page limitations her, please refer to our earlier paper [2]. Additionally, [2] provides more information on how the random PV output for each agent is generated. The power flow equations (7) are solved by the open-source distribution system simulator OpenDSS [34].
5.2 Deep Neural Network Implementation
@@
5.3 Comaprison of Three Different MARL Algorithms
@@
$\widehat{\nabla_{\theta_{i}}J_{i}\left(\theta_{i}\right)}=\mathbb{E}_{s\sim\rho^{\theta},a_{i}\sim\pi_{\theta_{i}}}\left[\nabla_{\theta_{i}}\log\pi_{\theta_{i}}\left(a_{i}\mid s_{i}\right)Q_{i}^{\pi}\left(\bm{s};\bm{a_{1},\ldots,a_{I}}\right)\right],$ (27)
@@
While the action-value function $Q_{i}^{\pi}$ is the same as in (27), the key difference lies in the policy function. As stated in Section 4.1 and reflected in (29), each agent $i$’s policy depends on all agents’ states $\bm{s}$. The other difference between the consensus MARL and the MADDPG algorithm is how the action-value function $Q_{i}^{\pi}$ is updated. In MADDPG, the function is updated through centralized training by a central authority; while in consensus MARL, updates are executed via a decentralized consensus process, as described in Section 4.2.
5.4 Numerical Results
We conduct 20 simulation runs for each of the three frameworks, each comprising 7,000 training episodes, where each training episode is 24 time steps (corresponding to 24 hours in a day). In Figure 6, the mean of the convergence curves of the 30-episode moving average of the episodic total reward in the distribution network is depicted for the 20 runs of each of the three distinct MARL algorithms. The shaded region represents one standard deviation.
@@ 5.4 Numerical Results
It is evident that the mean episodic total reward gradually converges to a high level at the end of the training in all three frameworks. The results show that the consensus MARL algorithm outperforms the fully decentralized framework, which is in line with expectations, since the consensus-update framework benefits from improved approximation of the action-value function through communication. A one-sided Welch’s t-test yields a p-value of $7.46\times 10^{-7}$, strongly suggesting the rejection of the null hypothesis that the episodic total reward of consensus MARL after 7,000 training episodes is less than or equal to that of MADDPG. This result provides strong evidence favoring consensus MARL for better performance compared to MADDPG, which is somewhat surprising. Conventionally, one might anticipate a centralized algorithm to outperform decentralized algorithms. The explanation here can be traced to MADDPG’s decentralized execution. In MADDPG, each agent independently updates its actions based on its own local state, despite the centralized nature of the learning process. This approach of relying solely on local information appears to be less effective compared to the consensus update mechanism.
Figure 7 presents the average market prices of the P2P market in chronological order over the last three days after thousands of training episodes.
@@
Notably, in the consensus-update framework, the market price tends to be higher during off-peak hours and lower during peak hours. This suggests that agents using this algorithm have developed more sophisticated trading strategies compared to the other two MARL frameworks. Specifically, the consensus-update algorithm leads to strategies where agents buy more and sell less during off-peak hours, and sell more and buy less during peak hours, unlike the purely decentralized algorithm.
Figure 8 shows the gradual reduction of the system voltage deviation after training. Initially, it is observed that voltage violations are significantly high across all three frameworks. [^4] We would like to emphasize that the violations depicted in the figures are not actual violations. They occur in the context of a preliminary stage, such as the hour-ahead market. If agents’ bids result in physical network constraint violations, these bids are returned to the agents for them to resubmit till the bids are feasible. In real-time operations, deviations from the cleared bids are inevitable. In such cases, we assume the DSO, typically a utility company, would have other resources available to ensure grid reliability and stability. The costs associated with maintaining safety and stability could be distributed among all market participants, similar to how T&D costs are allocated, as mentioned earlier. However, the consensus and MADDPG framework demonstrate a rapid decline in these violations, converging swiftly and effectively to zero at the fastest rate at the later stage. The fully decentralized framework demonstrates a slower rate of convergence, and is seemingly unable to fully eliminate voltage violations.
@@
Figure 9: Transmission line capacity violation. The transmission line capacity violation (kVA) in the $n$-th episode is calculated as $\sum_{t=24(n-1)}^{24n-1}\max\{0,S^{684,671}_{t}-15\}$.
Figure 9 illustrates the effectiveness of the three MARL frameworks in reducing distribution line capacity violations. However, the decrease is not always consistent, especially at the beginning of each training cycle for MADDPG. This initial fluctuation can be attributed to the core objective of all MARL frameworks: optimizing the total long-term reward through a balance between exploration (trying new actions) and exploitation (focusing on actions with proven success). During training, certain components of the total reward might temporarily prioritize actions that lead to short-term capacity violations, causing these fluctuations in the observed reduction.
6 Conclusion and Future Research
This study has developed a market and algorithmic framework to enable consumers’ and prosumers’ participation in local P2P energy trading. Utilizing reinforcement learning algorithms, we have automated bidding for agents while ensuring decentralized decision-making. The SDR-based market clearing addresses challenges with zero-marginal-cost resources and simplifies bidding. Additionally, our MARL framework includes voltage and line capacity constraints of a physical network, setting a foundation for real-world applications.
@@ 6 Conclusion and Future Research
Acknowledgment
We acknowledge the support of the U.S. Department of Energy, Office of Electricity under Award Number DE-OE0000921, and the National Science Foundation under Grant No. ECCS-2129631.